EXPERIMENTAL EVIDENCE OF LONE PAIRS - π SYSTEM INTERACTION: THE ROTATIONAL SPEC-TRUM OF CHLOROTRIFLUOROETHYLENE - WATER COMPLEX

\underline{L. EVANGELISTI}, Q. GOU, G. FENG and W. CAMINATI, Dipartimento di Chimica ”G. Ciamician” dell’Universitá, Via Selmi 2, I-40126 Bologna, Italy.

Chemists have always been interested in labeling the interactions between the molecules, i.e. from covalent to ionic bond or van der Waals force. Certainly the most important non-covalent bond is the hydrogen bond, also of fundamental importance in biology. More recently also weak Hbond, sometimes in competion with halogen bonding, have been investigated because they have assumed a fundamental importance. Here, we show the lone pairs - π interaction prevails on the latter ones. We measured the molecular beam Fourier transform microwave spectra of five isotopologues of the 1:1 adduct of chlorotrifluoroethylene with water. Besides the rotational constants, the quadrupole coupling constants of the chlorine atom have been determined. Quantum chemistry calculations, at the MP2/6-311++G(d,p) level, have been carried out in order to obtain information about the structure and relative stability of the conformers under study.